OpenStudy (anonymous):

Passing through (-2,1) and parallel to the line whose equation is x-3y=7. write an equation for the line in slope intercept form. Thats all I need because i keep messing up

OpenStudy (anonymous):

You should first write the equation of the known line into slope-intercept form. You'll then have to parameters: the slope and the offset (or intercept). Which one of these matters when talking about parallel lines? Once you have figured that out, you can use the point to determine the other parameter.

OpenStudy (anonymous):

req eq will be (x+2)=1/3(y-1) so 3x+6-y+1=0 so 3x-y+7=0 so y=3x+7 is ans

OpenStudy (anonymous):

I don't think that is the right answer, parik. y=3x+7 does go through (-2,1), but it is not parallel to the other line. Your line has a slope of 3, while the other line has a slope of 1/3.

OpenStudy (anonymous):

begin with x-3y=7 solve for y, to give slope intercept form: x-7=3y therefore y=x/3 -7/3 and m (slope)= 1/3 substitute the slope into a point slope form \[y-y_1 = m(x-x_1)\] where \[y_1 \] and \[x_1\] are the Cartesian (x.y) coordinates. Therefore, your equation should read:\[y-1=1/3 (x+2)\] then simply solve for y again: \[y=x/3 + 5/3\] and that will be your answer